# LLMEasyQuant - An Easy to Use Toolkit for LLM Quantization

### v1

@@ Abstract
1University of Wisconsin-Madison,
2University of Notre Dame
@@ 1 Introduction
Traditional quantization methods like uniform and scalar quantization, emphasizing their utility in diverse applications such as telecommunications and consumer electronics.
@@ 2 Methodology
LLMEasyQuant is a toolkit designed to simplify the process of quantizing large language models (LLMs). To achieve an easy quantization method, we developed LLMEasyQuant as an easy-to-use toolkit for straightforward quantization.
@@
  • smoothquant: A smoothing technique to the quantization process.[XLS+24]
  • symquant: Symmetric scaling based on the absolute maximum value.[YAZ+22]
@@ 3.1.2 Mathematical Formulation
3.1.3 Applications
ZeroQuant is commonly used in areas where data sparsity is significant, such as in the storage and transmission of high-dimensional but sparse datasets. In neural networks, employing ZeroQuant can significantly reduce the computational costs and power consumption during inference on edge devices.
@@ 3.2.2 Mathematical Formulation
3.2.3 Applications
This quantization method is extensively used in hardware implementations of neural networks, where maintaining data symmetry simplifies the computational requirements and enhances performance consistency.
@@ 3.3.3 Implementation Example
3.3.4 Applications
Particularly beneficial in deploying deep neural networks on mobile or embedded devices where memory and processing power are limited.
@@ 3.5.5 Dequantization Process
3.6 Applications
SimQuant is particularly effective in environments where storage and processing resources are at a premium, such as in embedded systems, mobile devices, and cloud-based machine learning platforms. Its ability to adaptively quantize data makes it suitable for real-time applications that require efficient data processing on-the-fly.
@@ 3.7.1 Definition and Concept
SmoothQuant [XLS+24] involves a sophisticated quantization approach that applies a smoothing algorithm to data before quantization. This method is particularly effective in applications where preserving the relational dynamics between different features of data is crucial.
@@ 3.7.3 Implementation Details
3.7.4 Applications
SmoothQuant is particularly beneficial in neural network training and inference, where it can lead to more stable and robust models by mitigating the impact of quantization noise. It’s also used in image processing and audio signal processing to maintain quality while reducing data bandwidth.

### v3

# LLMEasyQuant: A User-Friendly Toolkit for LLM Quantization

@@ 1 Introduction
With the development of Large Language Models (LLMs), the models have grown extremely large, so the memory usage and inference speed are greatly limited by the size of the model. Consequently, as one of the most popular technique for model compression, quantization has many variants now used for LLM compression and inference acceleration. The goal of quantization in LLMs is to reduce their size while minimizing its influence on inference speed.
@@ 2 Methodology
LLMEasyQuant is a toolkit designed to simplify the process of deploying quantization on large language models (LLMs). Its functions are included as follows:
@@
  • symquant: Symmetric quantization, a quantization method for scaling based on the absolute maximum value.[FFBL18]
  • smoothquant: A quantization method to achieve efficient post-training quantizatio for LLMs.[XLS+24]
@@ 3.5.5 Dequantization Process
3.5.6 Pseudocode for SimQuant
Algorithm 1 KVQuant Quantization
1:function QuantizeKVCache(Key, Value, scaling_factor, zero_point)
2:     for each layer $l$ do
3:         $Key_{l}\leftarrow\text{Quantize}(Key_{l},scaling\_factor,zero\_point)$
4:         $Value_{l}\leftarrow\text{Quantize}(Value_{l},scaling\_factor,zero\_point)$
5:     end for
6:     return $Key,Value$
7:end function
@@ 3.6.1 Definition and Concept
SmoothQuant [XLS+24] is a post-training quantization (PTQ) method designed to facilitate accurate and efficient quantization of large language models (LLMs). It aims to achieve 8-bit quantization for both weights and activations without retraining, ensuring minimal loss in model accuracy. SmoothQuant addresses the challenge of activation quantization by redistributing the quantization difficulty from activations to weights, which are inherently easier to quantize.
The key concept behind SmoothQuant is the smoothing of activation outliers, which typically hinder effective quantization. By migrating the quantization difficulty from activations to weights, SmoothQuant ensures that both can be quantized to 8-bit integers (INT8) effectively. This migration is achieved through a mathematically equivalent transformation that adjusts the scale of activations and weights.
@@ 3.6.2 Mathematical Formulation
3.6.3 SmoothQuant Methodology
SmoothQuant smooths activation outliers by migrating quantization difficulty from activations to weights. The following pseudocode outlines the main steps of the SmoothQuant process.
Algorithm 2 SmoothQuant Process
1:Input: Activation tensor $X$, Weight tensor $W$
2:Output: Quantized activation $\overline{X}$, Quantized weight $\overline{W}$
3:Compute per-channel scaling factor $s_{j}=\max(|X_{j}|)^{\alpha}/\max(|W_{j}|)^{1-\alpha}$
4:for each channel $j$ do
5:     Smooth activation $X_{j}\leftarrow X_{j}/s_{j}$
6:     Adjust weights $W_{j}\leftarrow W_{j}\times s_{j}$
7:end for
8:Quantize smoothed activation and adjusted weights
[…]
11:return $\overline{X},\overline{W}$
@@ 5 Discussions
LLMEasyQuant combined with downstream applications such as [WLC+24], [LZC24], [LWC+24] can achieve satisfactory performance.